OpenStudy (anonymous):

A particle starts at the origin and moves along the curve y=2x3/2 in the positive x-direction at a speed of 3 cm/sec, where x,y are in cm. find the position of the particle at t=6. Now the answers are x = 8.221 and y = 15.714 but how you get there remains a mystery to me. Thanks for helping

OpenStudy (anonymous):

Slight edit it's y = (2/3)*x^(3/2)

OpenStudy (anonymous):

Since it's talking about three variables: x, y, and t. Perhaps parametric equations?

OpenStudy (anonymous):

Hmm, hold on.. does it mean the speed along the curve is 3cm/s or just the speed in the X direction is 3cm/s?

OpenStudy (anonymous):

To be honest I'm not quite sure, these questions are very poorly written and this is exactly how it is in the book

OpenStudy (anonymous):

I tried it with the assumption that the speed was in the X direction only and didn't get those numbers you listed, so I guess it must be the other way. (that complicates things . . .)

OpenStudy (anonymous):

I've been working on this one forever and I just don't know how to get to those results

OpenStudy (anonymous):

You know what I mean by parametric functions, though, right? Need to find x=f(t) and y=g(t)=(2/3)f(t)^(3/2).

OpenStudy (anonymous):

oh yeah of course

OpenStudy (anonymous):

If it's speed along the curve, then makes me think that finding arc length will be necessary.

OpenStudy (anonymous):

well arc length would be integral from a to b for speed

OpenStudy (anonymous):

and speed is sqrt((x')^2 + (y')^2) or sqrt(1+ t) or at least that's what i'm getting

OpenStudy (anonymous):

If that x' and y' are derivatives with respect to t, then yes, I think that's correct, but we don't have x and y as functions of t yet. This sounds like a big pain in the neck, but I think this strategy might work: \[arc \space length, \space s=\int\limits_{0}^{x} \sqrt{1+(\frac{dy}{dx})^2} dx\] Distance = rate × time, and the rate is 3cm/s for 6s, so s=18cm. Set that integral equal to 18 and solve for the upper limit, x.

OpenStudy (anonymous):

well we can just say x = t and y = (2/3)t^(3/2)

OpenStudy (anonymous):

I'm going to toss up the Bat-Signal. @TuringTest @ash2326 @experimentX @hartnn Second opinion on this, guys?

OpenStudy (anonymous):

We don't know if x=t is the proper function. I tried x=3t when I assumed that the speed was only in the X direction at first, but that didn't work.

OpenStudy (anonymous):

If x=t then the position at t=6 would be (6, 9.797..)

OpenStudy (anonymous):

The arc length formula isn't too bad here because (dy/dx)^2 = x. The integral isn't very messy.

OpenStudy (anonymous):

Dang, I think I'm missing something. I tried it and got exactly 8 for x... So close . . .

OpenStudy (experimentx):